# Exploring Next Token Prediction For Optimizing Databases

### v2

@@ 4. The Probe and Learn (PoLe) Framework
To effectively adopt the Next Token Prediction (NTP) paradigm in database systems, we propose a novel technique termed Probe and Learn (PoLe, for short). For a given DBMS optimization task, this involves probing the hardware during query execution to generate DB-tokens. These tokens are used to train a DT to learn an optimal policy for the specified optimization task. Refer to Figure 1 for a detailed overview of the PoLe technique.
Probe: The Probe step comprises of the following 4 steps.
@@
  • Choosing. For a given optimization task with a set objective, i.e., query throughput, tail latency or scalability, we execute different policies across various hardware configurations and workloads, drawn from a diverse set of policy, hardware, and workload pools. This includes both optimal and suboptimal policies.
  • Profiling. During query execution, we periodically profile the hardware to capture the behavior of crucial hardware components using hardware statistics, e.g., the number of instructions executed, L1-d, L1-i, Last Level Cache (LLC, for short), and branch misses.
@@
  • Tokenization. Alongside the hardware profiles, we monitor the desired performance metric of the database system. These hardware profiles along with the performance metric serve as the DB-tokens. These DB-tokens are analogous to the state and RTG-tokens in DT. We also tokenize the DBMS policy to generate action tokens.
@@
  • Offloading. We periodically offload these action tokens along with the associated DB-tokens to an offline dataset for the learning phase.
Learn: The Learn step comprises of the following 2 steps.
@@
  • Training. We train a Decision Transformer (DT) on the collected offline dataset following the Teacher Forcing strategy (Williams and Zipser, 1989). As the DT is trained on diverse policies across varying hardware and workload configurations, it builds a mental model of the performance landscape of the optimization task in hand. It gains insights into the inherent tradeoffs and the performance improvements across different strategies, hardware and workload via the interaction of action and DB-tokens, that are present in the offline dataset.
@@
  • Inference. Using the trained DT, we infer a new policy auto-regressively. The process begins by feeding DT with the initial DB-token, i.e., state and the desired objective of the specified optimization task. In turn, DT generates the next action token which is fed back to the DBMS to generate the subsequent DB-token. The DB-token is then fed back into DT to get the next action token. This iterative process continues until the complete policy is generated. Note that, during the inference process, the DT has no explicit knowledge about the hardware that the DBMS is running on, or the workload that the DBMS is executing. Hence, both the hardware, and/or the workload can be unseen. It only relies on the DB-tokens, and the inherent knowledge it has gained by being trained on a wide range of policies, hardware configurations, and workloads to generate the next action token.
@@
Case Study: Scheduling Indexing Tasks Over NUMA Servers. To demonstrate the effectiveness of PoLe , we apply it to the main-memory index scheduling task. Modern Non Uniform Memory Access (NUMA, for short) servers are notoriously complex and heterogeneous. Each server can contain more than 1000 cores (Anandtech, 2024) spread across multiple processor dies. Data access latencies can vary significantly by up to 4$\times$ depending on the distance between the core that initiates the data request and the DIMM chip where the data resides. Additionally, concurrently running queries can cause interference by congesting on-chip and off-chip interconnects, as well as saturating hardware queues at caches and memory controllers. Thus, the performance of a DBMS index can vary significantly depending on the physical location of the core, where the queries are scheduled, and the DRAM chips where data resides.
@@
We formulate main-memory index scheduling as an NTP task and adopt the PoLe framework to solve it. The index is divided into multiple ($\mathcal{I}_{c}$) index chunks, i.e., a set of index nodes. Given a set of compute cores, the goal is to predict the core on which queries should be scheduled for the next index chunk. The data placement is implicit, i.e., the data for each chunk resides in the DIMM chip associated with the assigned core. The core ids serve as the action tokens. The clock-cycles, cache, branch misses, local and remote DRAM accesses along with the query throughput constitute the DB-token. A complete policy includes a sequence of $\mathcal{I}_{c}$ core ids. The $i$th core in the sequence represents the core assigned to process the corresponding index slice.
5. Initial Evaluation
@@
We evaluate the learned scheduling policy by PoLe against various OS policies and heuristics. The default OS policy (OS:D) is local allocation, i.e., the OS kernel places an index node on the NUMA node of the requesting CPU. The OS:I policy places the index nodes in an interleaved manner across all the NUMA nodes. For both the Shared Everything NUMA (SE:N) and the Shared Nothing (Porobic et al., 2012) (SN) strategies, nearby index slices are clustered together and placed in the same NUMA node. For OS:D, OS:I, and SE:N, the OS handles the scheduling. For SN, the scheduling strictly follows the data placement. We run a YCSB-like benchmark (Cooper et al., 2010) on a main-memory B+-Tree on the YCSB-A workload. Initially, the index is loaded with 30M records. Each index record has a 64-bit key and a 64-bit value. The DT in the experiment is trained on tokens generated by executing different YCSB workloads, e.g., YCSB-A, YCSB-C, YCSB-E, on a wide range of hardware configuration. Figures 2a, 2b show that PoLe can learn better scheduling policies on Intel Sandy Bridge and NVIDIA Grace Hopper, outperforming the baselines by up to 2.78$\times$. For both these experiments, the workload and the hardware is seen, i.e., the offline dataset that DT has been trained on contains experiences from the same hardware and the same workload. Figure 2c shows that PoLe can learn a better scheduling policy even for an unseen hardware, i.e., the offline dataset that DT is trained on does not contain any experience for the particular hardware. It outperforms the baselines by 3$\times$ on an unseen Intel Skylake X 4 Socket 4 NUMA Node machine. Thus, the NTP paradigm not only enhances performance on existing workloads and hardware but also generalizes to optimize index performance across previously unseen hardware.
@@ 6. Concluding Remarks
In this paper, we lay the foundation for incorporating the Next Token Paradigm in optimizing database systems by proposing a framework that generates hardware tokens to train Decision Transformers in a data-driven framework. Whether all or which database optimization tasks can benefit from the proposed PoLe framework as a means to generalize across diverse hardware and workloads is yet to be seen. However, preliminary results on index scheduling over NUMA servers suggest that PoLe has strong potential in improving performance and adapting to diverse environments.

### v3

@@ 4. The Probe-and-Learn (PoLe) Framework
To effectively adopt the Next Token Prediction (NTP) paradigm in database systems, we propose a new technique, termed Probe-and-Learn (PoLe, for short). For a given DBMS optimization task, this involves probing the hardware during query execution to generate DB-tokens. These tokens are used to train a DT to learn an optimal policy for the specific optimization task. Refer to Figure 2 for an overview of PoLe.
Figure 3. Performance of a main-memory B+-Tree under various scheduling policies.
Probe: The Probe step comprises the following four steps.
@@
  • Choosing. For a given database optimization task with a set objective, i.e., query throughput, tail latency or scalability, we execute different policies across various hardware configurations and workloads, drawn from a diverse set of policy, hardware, and workload pools. The policy pool includes both optimal and suboptimal strategies that can be randomly sampled or borrowed from existing database literature. The workload pool can comprise a mixture of transactional and analytical workloads. The hardware pool can span servers from different vendors, with varying micro-architectural properties, NUMA configurations, memory technologies, and other distinct characteristics, all of which can significantly impact the performance and behavior of the DBMS.
  • Profiling. During query execution, we periodically profile the hardware to capture the behavior of crucial hardware components using hardware statistics, e.g., the number of instructions executed, L1-d, L1-i, L2, Last Level Cache (LLC, for short) accesses and misses, branch mispredictions, the number of local and remote memory accesses and misses, etc. Depending on the DBMS optimization task, the granularity of profiling can vary. For example, for scheduling, profiling can be performed at the query level. In contrast, the join order selection and the data partitioning task requires profiling at the level of query operators and data partitions, respectively. Similarly, different optimization tasks can focus on different hardware statistics. For example, NUMA-aware optimizations can focus on local and remote memory accesses to capture the data locality. In contrast, the join order selection task can benefit from the write statistics as they reflect the size of the materialized intermediate results.
@@
  • Tokenization. Alongside the hardware profiles, we monitor the desired performance metric of the database system, e.g., the average query throughput, the average tail latency, scalability, etc. These hardware profiles along with the performance metric serve as the DB-tokens. These DB-tokens are analogous to the state and RTG-tokens in DT. Moreover, we tokenize the DBMS policies to generate action tokens.
@@
  • Offloading. Periodically, we offload the action tokens along with the associated DB-tokens to an offline dataset to minimize cache and memory pollution, incurred by this additional data. The offline dataset is provided to the DT during the learning phase.
Learn: The Learn step comprises the following two steps.
@@
  • Training. We train a Decision Transformer (DT) on the collected offline dataset in a supervised manner following the Teacher Forcing strategy (Williams and Zipser, 1989). As the DT is trained on diverse policies across varying hardware and workload configurations, it builds a mental model of the performance landscape of the optimization task at hand. It gains insights into the inherent tradeoffs and the performance improvements across different strategies, hardware and workload via the interaction of action and DB-tokens, that are present in the offline dataset.
@@
  • Inference. Using the trained DT, PoLe infers a new policy using auto-regression. The process begins by feeding DT with the initial DB-token, i.e., state and the desired objective of the specified optimization task. In turn, DT generates the next action token which is fed back to the DBMS to generate the subsequent DB-token. The DB-token is then fed back into DT to get the next action token. This iterative process continues until the complete policy is generated. Note that, during the inference process, the DT has no explicit knowledge about the hardware that the DBMS is running on, or the workload that the DBMS is executing. Hence, both the hardware, and/or the workload can be unseen. It only relies on the DB-tokens, and the inherent knowledge it has gained by being trained on a wide range of policies, hardware configurations, and workloads to generate the next action token.
5. Case Study: Scheduling Indexing Tasks Over NUMA Servers
@@
To demonstrate the effectiveness of PoLe, we apply it to the main-memory index scheduling task. Modern Non Uniform Memory Access (NUMA, for short) servers are notoriously complex and heterogeneous. Each server can contain more than 1000 cores (Anandtech, 2024) spread across multiple processor dies. Data access latencies can vary significantly by up to 4$\times$ depending on the distance between the core that initiates the data request and the DIMM chip where the data resides. Additionally, concurrently running queries can cause interference by congesting on-chip and off-chip interconnects, as well as saturating hardware queues at caches and memory controllers. Thus, the performance of a DBMS index can vary significantly depending on the physical location of the core where the queries are scheduled, and the DRAM chips where the corresponding data resides.
@@
We formulate main-memory index scheduling as an NTP task and adopt the PoLe framework to solve it. The index is divided into multiple ($\mathcal{I}_{c}$) index chunks, i.e., a set of index nodes. Given a set of compute cores, the goal is to predict the core on which the queries should be scheduled for the next index chunk. The data placement is implicit, i.e., the data for each chunk resides in the DIMM chip associated with the assigned core. The core ids serve as the action tokens. The clock-cycles, cache, branch misses, local, and remote DRAM accesses along with the query throughput constitute the DB-token. A complete policy includes a sequence of $\mathcal{I}_{c}$ core ids. The $i$th core in the sequence represents the core assigned to process the corresponding index slice.
Figure 4. The scheduling policies learned by PoLe. Each grid corresponds to the learned scheduling policy for the corresponding server for the YCSB-A workload. The top-left grid cell corresponds to the 1st index chunk. The first cell in $i$-th row corresponds to the $\langle i*16+1\rangle$-th index chunk. The corresponding text of a grid cell signifies the hardware core where the query for that particular index chunk is scheduled. Cells sharing the same color indicate that the associated index chunks are scheduled on the same NUMA server.
Figure 5. Top performing scheduling policies present in the training dataset for the Intel Sandy Bridge and NVIDIA Grace Hopper servers. Each grid corresponds to a scheduling policy for the corresponding server. The same scheduling policy can be present multiple times in the training dataset for different workloads. Cells sharing the same color indicate that the associated index chunks are scheduled on the same NUMA server.
6. Evaluation and Preliminary Results
Experiment settings: We evaluate the learned scheduling policy by PoLe against various OS policies and heuristics. The default OS policy (OS:D) is local allocation, i.e., the OS kernel places an index node on the NUMA node of the requesting CPU. The OS:I policy places the index nodes in an interleaved manner across all the NUMA nodes. For both the Shared Everything NUMA (SE:N) and the Shared Nothing (Porobic et al., 2012) (SN) strategies, nearby index slices are clustered together and are placed in the same NUMA node. For OS:D, OS:I, and SE:N, the OS handles the scheduling. For SN, the scheduling strictly follows the data placement.
We run a YCSB-like benchmark (Cooper et al., 2010) on a main-memory B+-Tree (Wang et al., 2018; Wang, 2017) on the YCSB-A workload for 5 minutes. Initially, the index is loaded with 30M records, and grows up to 200M records as more insertions are performed. Each index record has a 64-bit key and a 64-bit value. For these experiments, the B+-Tree is divided into 256 index chunks. The scheduling policy involves allocating the CPU cores for these 256 index chunks.
We evaluate the scheduling policies on the following three machines: Intel Sandy Bridge, NVIDIA Grace Hopper, and Intel Skylake X (Unseen Hardware). The Intel Sandy Bridge is a 4 socket, 4 NUMA Node, 32 ($\times$2) core machine running Ubuntu 22.04 on a CloudLab (Duplyakin et al., 2019) d820 node. Each socket in the machine is an 8 ($\times$2) core Intel Xeon E5-4620 processor, with an inclusive 64MB L3 cache. Each physical core in the hardware has a 1MB L1-d, 1MB L1-i, and an 8MB L2 cache. The machine is equipped with 128 GB of DDR3 RAM, and the CPU clock frequency is 2.2 GHz. It does not support Sub-NUMA clustering (SNC) or Cluster On Die (COD) technology. The NVIDIA Grace Hopper is a single socket 72 core machine running Ubuntu 22.04 on a CloudLab (Duplyakin et al., 2019) nvidiagh node. The processor is based on an ARM’s Neoverse V2 architecture, where the CPU frequency is 3.1 GHz. The machine is equipped with 480GB LPDDR5X RAM. The Skylake-X is a 4 socket, 4 NUMA Node, 92 (×2) core machine running on Ubuntu 22.04. The respective size of L1-d, L1-i, L2, and LLC cache are 3MB, 3MB, 96MB, and 132MB, respectively. The CPU clock frequency is 2.7 GHz.
@@ 6. Evaluation and Preliminary Results
Preliminary Experimental Results. The DT in the experiment is trained on tokens generated by executing different YCSB workloads, e.g., YCSB-A, YCSB-C, YCSB-E, on a wide range of hardware configurations. Figures 3a and 3b illustrate that PoLe can learn better scheduling policies on Intel Sandy Bridge and NVIDIA Grace Hopper, outperforming the baselines by up to 2.78$\times$. For both these experiments, the workload and the hardware have been encountered during training, i.e., the offline dataset that the DT has been trained on contains experiences from the same hardware and the same workload. Figure 3c illustrates that PoLe can learn a better scheduling policy even for an unseen hardware, i.e., the offline dataset that DT has been trained on does not contain any experience for the particular hardware. It outperforms the baselines by 3$\times$ on an unseen Intel Skylake X 4-Socket 4-NUMA-Node machine. Thus, the NTP paradigm not only enhances performance on existing workloads and hardware but also generalizes to optimize index performance in unseen hardware that was not encountered during the training phase.
Figure 4 gives the scheduling policies learned by the DT for the three corresponding servers. Each grid cell corresponds to an index chunk. The number associated with a grid cell corresponds to the hardware core where the query for that particular index chunk has been scheduled. To ensure that the DT does not merely replicate the best scheduling policies from the offline training dataset, we present the top scheduling policies observed during training in Figure 5, and it confirms that the DT does not replicate the scheduling policies from the offline training dataset.
Offline Dataset Preparation Cost: The offline training dataset is generated by executing a diverse set of scheduling policies on a wide range of YCSB workloads on an Intel Sandy Bridge, and an NVIDIA Grace Hopper server. Precisely, the offline dataset contains 445 and 49 scheduling episodes, i.e., samples on the two respective servers. Each episode lasts around 5 minutes. Renting an Intel Sandy Bridge-equivalent Haswell server, i.e., a x1.16xlarge (Amazon, 2025b) instance on AWS costs approximately \$614 (\$1.380/hr), while renting an NVIDIA Grace Hopper server on Lambda cloud (Lambda, 2025) costs approximately \$73 (\$1.49/hr), totaling an approximate cost of \$687.111The AWS pricing info are collected from: https://instances.vantage.sh/.
Training Cost: We train PoLe’s DT on a dual-socket 64 (×2) core machine running Ubuntu 22.04 on a CloudLab (Duplyakin et al., 2019) r6525 node. The DT comprises approximately 3.37M trainable parameters. It takes approximately 26 CPU hours to train the DT to reach a training accuracy of over 95%. The inference takes up to 1 CPU minute on the same machine. Renting an equivalent m5a.24xlarge (Amazon, 2025a) Amazon AWS instance costs approximately \$32 (\$1.211/hr) to train the DT used in the experiment.
@@ 7. Concluding Remarks
In this paper, we lay the foundation for incorporating the Next Token Paradigm in optimizing database systems by proposing a framework that generates hardware tokens to train Decision Transformers in a data-driven framework. Whether all or which database optimization tasks can benefit from the proposed PoLe framework as a means to generalize across diverse hardware and workloads is yet to be seen. However, preliminary results on index scheduling over NUMA servers suggest that PoLe has strong potential in improving performance and on adapting to diverse environments. Building on these early promising results, a key direction for future work is to conduct an extensive investigation into the database optimization landscape to categorize the optimization tasks that can benefit from the Next Token Prediction (NTP) paradigm, and assess to what extent the PoLe framework can provide consistent performance and adaptivity guarantees.